b1az3:

Solve the following system. y = -2x 2 + 3x - 1 and x + y = -1. The solutions are and

jhonyy9:

rewrite them in this way and take them equal y = -2x^2 +3x -1 y = -1 -x -2x^2 +3x -1 = -1 -x solve this quadratic for x then using these x_1,2 you can calcule easy the y values hope helped understandably easy

b1az3:

2,0

jhonyy9:

-2x^2 +3x -1 = -1 -x add to both sides 1 and get -2x^2 +3x = -x add to both sides x and get -2x^2 +4x = 0 factorize out the -2x and get -2x(x -2) = 0 -2x = 0 x_1 = 0 x-2=0 x_2 = 2 and the y_1,2 values ?

jhonyy9:

x+y = -1 use this equation to calcule y_1,2 = ?

jhonyy9:

@AZ